# A causal inference framework for spatial confounding

### v7

@@ 1 Introduction
For most of its history, the field of spatial statistics has been concerned primarily with spatial process models, where the goal is prediction and interpolation of a spatially varying outcome $Y$ (Banerjee et al., 2008; Cressie and Wikle, 2011). However, there is now increasing interest in making inferences about the causal effect of an exposure $X$, possibly also spatially varying, on $Y$. Estimation of such causal effects needs to mitigate potential spatial confounding: unmeasured variables that are thought to possess a spatial structure and influence both the exposure and outcome of interest. Researchers have attempted to clarify when existing spatial models can support valid causal inferences (Paciorek, 2010; Hodges and Reich, 2010; Schnell and Papadogeorgou, 2020; Keller and Szpiro, 2020). These conclusions have often relied on a formalism of spatial confounding that is tied to the analysis model used. A few papers have drawn from causal inference theory and methods to propose adaptations to existing spatial models (Papadogeorgou et al., 2018; Schnell and Papadogeorgou, 2020; Guan et al., 2020; Osama et al., 2019; Thaden and Kneib, 2018; Dupont et al., 2020). These approaches still rely to some degree on parametric assumptions (e.g., linearity) used in the analysis model.
@@ 3.4 Shift interventions
[…]
@@ 7 Application
[…]
@@
[…]
@@
[…]
@@ 8 Discussion
Method
Interval estimation procedure
OLS / RSR (Hodges and Reich, 2010) (unadajusted estimator)
Bootstrap.
spline: PLM with spatial effect modeled by thin plate regression spline (Wood, 2003b). Implemented by mgcv (Wood, 2003a).
Analytic confidence intervals.
gp: PLM with spatial effect modeled by two-dimensional nearest-neighbors Gaussian process (NNGP) . Implemented by BRISC package (Saha and Datta, 2020).
BRISC bootstrap (Saha and Datta, 2018).
gSEM: geo-additive structural equation model (Thaden and Kneib, 2018). Splines fit by mgcv.
Bootstrap.
spatial+: spatial+ (Dupont et al., 2020). Splines fit by mgcv.
Bootstrap.
svc_mle: Spatially-varying coefficient model (Dambon et al., 2021a). Implemented by varycoef package (Dambon et al., 2021b).
Omitted (since resampled points cannot be accommodated by the GP covariance function).
spline_interaction: Spline with spatial interaction of exposure and space (hence, a spatially-varying coefficient model). Implemented by mgcv.
Bootstrap.
gp3: Three-dimensional Gaussian process over location and exposure (flexible model). Implemented by RobustGaSP (Gu et al., 2020).
Omitted (since resampled points cannot be accommodated by the GP covariance function).
spline3: Three-dimensional spline over location and exposure (flexible model). Implemented by mgcv.
Omitted due to computational constraints.
rf: Random forest over location and exposure (flexible model) (Breiman, 2001). Implemented by randomForest package (Liaw and Wiener, 2002).
Bootstrap.
BART: Bayesian additive regression tree (BART) over location and exposure (flexible model) (Chipman et al., 2010). Implemented by dbarts (Dorie, 2020) package.
Bootstrap.
DML_gp3: gp3 model augmented with spline model over space for exposure (flexible model, DML).
Omitted (since resampled points cannot be accommodated by the GP covariance function).
DML_spline: spline3 model augmented with spline model over space for exposure (flexible model, DML).
Omitted due to computational constraints.
DML_BART: BART model augmented with spline model over space for exposure (flexible model, DML).
Bootstrap.
@@
n method bias sd mse coverage
$1.000\times 10^{3}$ RSR $7.733\times 10^{-2}$ $1.603\times 10^{-2}$ $6.237\times 10^{-3}$ 0%
$1.000\times 10^{3}$ spline $-9.958\times 10^{-5}$ $6.297\times 10^{-3}$ $\mathbf{3.958\times 10^{-5}}$ 96%
$1.000\times 10^{3}$ gp $7.932\times 10^{-4}$ $6.509\times 10^{-3}$ $4.292\times 10^{-5}$ 95%
$1.000\times 10^{3}$ gSEM $1.362\times 10^{-3}$ $7.017\times 10^{-3}$ $5.099\times 10^{-5}$ 99%
$1.000\times 10^{3}$ spatial+ $5.348\times 10^{-4}$ $6.477\times 10^{-3}$ $4.216\times 10^{-5}$ 98%
$1.000\times 10^{3}$ svc_mle $3.294\times 10^{-4}$ $6.434\times 10^{-3}$ $4.143\times 10^{-5}$
$1.000\times 10^{3}$ spline_interaction $-9.566\times 10^{-5}$ $6.338\times 10^{-3}$ $4.010\times 10^{-5}$ 99%
$1.000\times 10^{3}$ gp3 $-5.282\times 10^{-4}$ $6.478\times 10^{-3}$ $4.216\times 10^{-5}$
$1.000\times 10^{3}$ spline3 $\mathbf{1.538\times 10^{-5}}$ $6.539\times 10^{-3}$ $4.268\times 10^{-5}$
$1.000\times 10^{3}$ rf $-1.886\times 10^{-1}$ $1.370\times 10^{-2}$ $3.575\times 10^{-2}$ 0%
$1.000\times 10^{3}$ BART $-2.754\times 10^{-3}$ $1.340\times 10^{-2}$ $1.868\times 10^{-4}$ 100%
$1.000\times 10^{3}$ DML_gp3 $-2.531\times 10^{-4}$ $6.690\times 10^{-3}$ $4.473\times 10^{-5}$
$1.000\times 10^{3}$ DML_spline $1.424\times 10^{-4}$ $6.749\times 10^{-3}$ $4.548\times 10^{-5}$
$1.000\times 10^{3}$ DML_BART $-1.811\times 10^{-3}$ $1.350\times 10^{-2}$ $1.850\times 10^{-4}$ 100%
$1.000\times 10^{4}$ RSR $7.715\times 10^{-2}$ $5.237\times 10^{-3}$ $5.980\times 10^{-3}$ 0%
$1.000\times 10^{4}$ spline $-2.203\times 10^{-4}$ $1.914\times 10^{-3}$ $3.696\times 10^{-6}$ 96%
$1.000\times 10^{4}$ gp $-1.842\times 10^{-4}$ $1.978\times 10^{-3}$ $3.932\times 10^{-6}$ 96%
$1.000\times 10^{4}$ gSEM $\mathbf{3.427\times 10^{-6}}$ $1.934\times 10^{-3}$ $3.724\times 10^{-6}$ 98%
$1.000\times 10^{4}$ spatial+ $-1.298\times 10^{-4}$ $1.912\times 10^{-3}$ $\mathbf{3.657\times 10^{-6}}$ 97%
$1.000\times 10^{4}$ spline_interaction $-2.171\times 10^{-4}$ $1.911\times 10^{-3}$ $3.684\times 10^{-6}$ 97%
$1.000\times 10^{4}$ spline3 $-1.965\times 10^{-4}$ $1.935\times 10^{-3}$ $3.768\times 10^{-6}$
$1.000\times 10^{4}$ rf $-9.426\times 10^{-2}$ $4.410\times 10^{-3}$ $8.904\times 10^{-3}$ 0%
$1.000\times 10^{4}$ BART $-8.981\times 10^{-4}$ $2.987\times 10^{-3}$ $9.694\times 10^{-6}$ 98%
$1.000\times 10^{4}$ DML_spline $-1.965\times 10^{-4}$ $1.991\times 10^{-3}$ $3.987\times 10^{-6}$
$1.000\times 10^{4}$ DML_BART $-3.668\times 10^{-4}$ $2.992\times 10^{-3}$ $9.053\times 10^{-6}$ 98%
@@
n method bias sd mse coverage
$1.000\times 10^{3}$ RSR $2.589\times 10^{-1}$ $1.528\times 10^{-2}$ $6.727\times 10^{-2}$ 0%
$1.000\times 10^{3}$ spline $6.863\times 10^{-3}$ $6.502\times 10^{-3}$ $8.930\times 10^{-5}$ 81%
$1.000\times 10^{3}$ gp $7.727\times 10^{-3}$ $6.770\times 10^{-3}$ $1.054\times 10^{-4}$ 76%
$1.000\times 10^{3}$ gSEM $7.305\times 10^{-2}$ $7.584\times 10^{-3}$ $5.394\times 10^{-3}$ 0%
$1.000\times 10^{3}$ spatial+ $1.084\times 10^{-2}$ $6.826\times 10^{-3}$ $1.641\times 10^{-4}$ 73%
$1.000\times 10^{3}$ svc_mle $7.387\times 10^{-3}$ $6.855\times 10^{-3}$ $1.015\times 10^{-4}$
$1.000\times 10^{3}$ spline_interaction $6.801\times 10^{-3}$ $6.568\times 10^{-3}$ $8.930\times 10^{-5}$ 96%
$1.000\times 10^{3}$ gp3 $2.682\times 10^{-3}$ $6.939\times 10^{-3}$ $5.524\times 10^{-5}$
$1.000\times 10^{3}$ spline3 $6.308\times 10^{-2}$ $4.068\times 10^{-2}$ $5.631\times 10^{-3}$
$1.000\times 10^{3}$ rf $-1.458\times 10^{-1}$ $1.595\times 10^{-2}$ $2.152\times 10^{-2}$ 0%
$1.000\times 10^{3}$ BART $7.672\times 10^{-3}$ $1.680\times 10^{-2}$ $3.405\times 10^{-4}$ 100%
$1.000\times 10^{3}$ DML_gp3 $\mathbf{1.396\times 10^{-3}}$ $7.065\times 10^{-3}$ $\mathbf{5.177\times 10^{-5}}$
$1.000\times 10^{3}$ DML_spline $2.998\times 10^{-2}$ $3.476\times 10^{-2}$ $2.104\times 10^{-3}$
$1.000\times 10^{3}$ DML_BART $3.432\times 10^{-3}$ $1.685\times 10^{-2}$ $2.950\times 10^{-4}$ 100%
$1.000\times 10^{4}$ RSR $2.580\times 10^{-1}$ $4.749\times 10^{-3}$ $6.661\times 10^{-2}$ 0%
$1.000\times 10^{4}$ spline $1.630\times 10^{-3}$ $1.947\times 10^{-3}$ $6.433\times 10^{-6}$ 86%
$1.000\times 10^{4}$ gp $2.159\times 10^{-3}$ $1.990\times 10^{-3}$ $8.605\times 10^{-6}$ 81%
$1.000\times 10^{4}$ gSEM $1.564\times 10^{-2}$ $2.089\times 10^{-3}$ $2.488\times 10^{-4}$ 0%
$1.000\times 10^{4}$ spatial+ $5.063\times 10^{-4}$ $1.958\times 10^{-3}$ $\mathbf{4.075\times 10^{-6}}$ 96%
$1.000\times 10^{4}$ spline_interaction $1.614\times 10^{-3}$ $1.945\times 10^{-3}$ $6.372\times 10^{-6}$ 88%
$1.000\times 10^{4}$ spline3 $2.244\times 10^{-2}$ $6.780\times 10^{-3}$ $5.495\times 10^{-4}$
$1.000\times 10^{4}$ rf $-7.677\times 10^{-2}$ $4.497\times 10^{-3}$ $5.914\times 10^{-3}$ 0%
$1.000\times 10^{4}$ BART $2.642\times 10^{-3}$ $3.618\times 10^{-3}$ $2.002\times 10^{-5}$ 97%
$1.000\times 10^{4}$ DML_spline $6.870\times 10^{-3}$ $7.080\times 10^{-3}$ $9.712\times 10^{-5}$
$1.000\times 10^{4}$ DML_BART $\mathbf{4.626\times 10^{-4}}$ $3.586\times 10^{-3}$ $1.302\times 10^{-5}$ 100%
@@
n method bias sd mse coverage
$1.000\times 10^{3}$ RSR $8.670\times 10^{-1}$ $1.173\times 10^{-1}$ $7.655\times 10^{-1}$ 0%
$1.000\times 10^{3}$ spline $1.070\times 10^{0}$ $7.224\times 10^{-2}$ $1.150\times 10^{0}$ 0%
$1.000\times 10^{3}$ gp $1.063\times 10^{0}$ $7.418\times 10^{-2}$ $1.135\times 10^{0}$ 0%
$1.000\times 10^{3}$ gSEM $1.123\times 10^{0}$ $7.587\times 10^{-2}$ $1.268\times 10^{0}$ 0%
$1.000\times 10^{3}$ spatial+ $1.090\times 10^{0}$ $7.215\times 10^{-2}$ $1.193\times 10^{0}$ 0%
$1.000\times 10^{3}$ svc_mle $2.019\times 10^{-2}$ $3.384\times 10^{-2}$ $1.551\times 10^{-3}$
$1.000\times 10^{3}$ spline_interaction $1.390\times 10^{-2}$ $3.400\times 10^{-2}$ $1.347\times 10^{-3}$ 94%
$1.000\times 10^{3}$ gp3 $3.863\times 10^{-3}$ $3.541\times 10^{-2}$ $1.266\times 10^{-3}$
$1.000\times 10^{3}$ spline3 $1.545\times 10^{-1}$ $1.074\times 10^{-1}$ $3.539\times 10^{-2}$
$1.000\times 10^{3}$ rf $-7.837\times 10^{-2}$ $4.106\times 10^{-2}$ $7.824\times 10^{-3}$ 67%
$1.000\times 10^{3}$ BART $1.024\times 10^{-1}$ $7.197\times 10^{-2}$ $1.566\times 10^{-2}$ 90%
$1.000\times 10^{3}$ DML_gp3 $\mathbf{3.363\times 10^{-3}}$ $3.541\times 10^{-2}$ $\mathbf{1.262\times 10^{-3}}$
$1.000\times 10^{3}$ DML_spline $1.439\times 10^{-1}$ $1.073\times 10^{-1}$ $3.220\times 10^{-2}$
$1.000\times 10^{3}$ DML_BART $9.417\times 10^{-2}$ $7.194\times 10^{-2}$ $1.403\times 10^{-2}$ 92%
$1.000\times 10^{4}$ RSR $8.687\times 10^{-1}$ $4.060\times 10^{-2}$ $7.562\times 10^{-1}$ 0%
$1.000\times 10^{4}$ spline $1.088\times 10^{0}$ $2.393\times 10^{-2}$ $1.183\times 10^{0}$ 0%
$1.000\times 10^{4}$ gp $1.085\times 10^{0}$ $2.416\times 10^{-2}$ $1.178\times 10^{0}$ 0%
$1.000\times 10^{4}$ gSEM $1.094\times 10^{0}$ $2.423\times 10^{-2}$ $1.197\times 10^{0}$ 0%
$1.000\times 10^{4}$ spatial+ $1.090\times 10^{0}$ $2.389\times 10^{-2}$ $1.189\times 10^{0}$ 0%
$1.000\times 10^{4}$ spline_interaction $\mathbf{2.216\times 10^{-3}}$ $1.080\times 10^{-2}$ $\mathbf{1.210\times 10^{-4}}$ 94%
$1.000\times 10^{4}$ spline3 $8.317\times 10^{-2}$ $2.091\times 10^{-2}$ $7.353\times 10^{-3}$
$1.000\times 10^{4}$ rf $-4.038\times 10^{-2}$ $1.200\times 10^{-2}$ $1.774\times 10^{-3}$ 16%
$1.000\times 10^{4}$ BART $3.066\times 10^{-2}$ $1.514\times 10^{-2}$ $1.168\times 10^{-3}$ 67%
$1.000\times 10^{4}$ DML_spline $6.955\times 10^{-2}$ $2.175\times 10^{-2}$ $5.308\times 10^{-3}$
$1.000\times 10^{4}$ DML_BART $2.832\times 10^{-2}$ $1.504\times 10^{-2}$ $1.027\times 10^{-3}$ 74%

### v8

@@ 1 Introduction
For most of its history, the field of spatial statistics has been concerned primarily with spatial process models, where the goal is prediction and interpolation of a spatially varying outcome $Y$ (Banerjee et al., 2008; Cressie and Wikle, 2011). However, there is now increasing interest in making inferences about the causal effect of an exposure $X$, possibly also spatially varying, on $Y$. Estimation of such causal effects needs to mitigate potential spatial confounding: unmeasured variables that are thought to possess a spatial structure and influence both the exposure and outcome of interest. Researchers have attempted to clarify when existing spatial models can support valid causal inferences (Paciorek, 2010; Hodges and Reich, 2010; Schnell and Papadogeorgou, 2020; Keller and Szpiro, 2020). These conclusions have often relied on a formalism of spatial confounding that is tied to the analysis model used. A few papers have drawn from causal inference theory and methods to propose adaptations to existing spatial models (Papadogeorgou et al., 2018; Schnell and Papadogeorgou, 2020; Guan et al., 2022; Osama et al., 2019; Thaden and Kneib, 2018; Dupont et al., 2020). These approaches still rely to some degree on parametric assumptions (e.g., linearity) used in the analysis model.
@@ 3.4 Shift interventions
The shift estimand is of direct policy relevance: it is more practical to consider interventions that incrementally change each individual’s exposure by some amount (for example, what would be the average effect on life-expectancy if PM2.5 levels were decreased everywhere by $2\mu g/m^{3}$?), rather than adjusting all individuals’ exposure levels to some common level $x$ which may be far from their current values. The former is captured by the shift intervention $E[Y(X)-Y(X-2)]$, where $X$ is PM2.5 with units $\mu g/m^{3}$, while the latter is the intervention directly considered by estimation of the exposure-response curve, or by a user-specified contrast $E[Y(x)-Y(x^{\prime})]$. Finally, by summarizing complex data-generating processes with a single scalar quantity, the shift estimand is valuable for its interpretability.
@@ 7 Application
Although the nature of the causal pathway is debated, birthweight has been widely studied for its strong correlation with health and economic outcomes in childhood and adulthood (Wilcox, 2001; Behrman and Rosenzweig, 2004; Chye and Lim, 1999); low birthweight is a risk factor for adverse outcomes including mortality. In addition, evidence suggests that exposure to ambient PM2.5 during pregnancy is linked to reductions in birthweight (Li et al., 2017; Basu et al., 2014). A 2021 meta-analysis gave an estimated $-27.55$ $g$ [95% CI: (-48.45, -6.65) ] change in birthweight for each $10\mu g/m^{3}$ additional concentration of exposure to PM2.5 during pregnancy, noting a wide range of previous effect estimates in the literature, all of which appear to derive from linear models (Uwak et al., 2021).
In order to estimate the causal effect of ambient PM2.5 on birthweight, it is necessary to adjust for confounding variables. “Greenspace” is a term with multiple definitions but generally refers to vegetation and natural spaces in residential environments (Taylor and Hochuli, 2017). The positive association between greenspace and birthweight has been studied (Toda et al., 2020; Cusack et al., 2017), as has the potential mitigation of PM2.5 by greenspace (Chen et al., 2019). Therefore, greenspace is a potential confounder of the relationship between PM2.5 exposure and birthweight. Since greenspace is a relatively smooth function of spatial location (i.e., nearby locations have similar levels of greenspace), we may attempt to control for greenspace by controlling for spatial location.
For the present study, we use records of births from the California Department of Public Health’s Birth Statistical Master Files with mother’s home addresses geocoded using ArcGIS; we restrict our analysis to conceptions in the year 2013. We obtain daily PM2.5 levels at 1 kilometer resolution from the NASA Socioeconomic Data and Applications Center (Di et al., 2021) and average them over the course of pregnancy. To measure greenspace, we use peak normalized difference vegetation index (NDVI) measured at 250-m resolution and aggregated at the census block level from the year 2013 (Didan et al., 2015). In all our analyses, we include the potential confounding variables: race/ethnicity, parity, Kotelchuck index, maternal age, maternal education, and census tract-level per capita income from the 2010 American Community Survey (sourced from Manson et al. (2023)). We exclude individuals with missing data, as well as pregnancies with recorded gestational lengths less than 22 weeks or greater than 44 weeks, mothers with recorded ages less than 15 or greater than 49, and birthweights recorded below 500 grams. This leaves us with $n=457,439$ live births. Latitude and longitude were recorded in the Web Mercator projection, and all quantitative variables were normalized prior to model-fitting.
(a) Average daily PM2.5 concentrations by census tract in 2013 (data from Centers for Disease Control and Prevention – National Environmental Public Health Tracking Network and Census Bureau (2010)). Note that the exposure values used in our study were more granular, as described in the text.
(b) Plot of average daily PM2.5 concentration exposure in $\mu g/m^{3}$ and birthweight in grams.
Figure 2: PM2.5 exposure effect study.
@@
Table 1: Estimated effects of a $0.1\mu g/m^{3}$ shift in average daily PM2.5 exposure on birthweight in grams. An asterisk indicates that the estimated smooth from the GAM model was used for adjustment in place of directly using the spatial location.
@@
Given that we fit multiple models, these results should be seen as exploratory rather than as a definite contribution to the scientific literature. Other limitations include the potential impact of “live-birth bias” (the bias arising from only including data from live births, analogous to the “birthweight paradox” concerning conditioning on low birth weight) (Goin et al., 2021; Hernandez-Diaz et al., 2006), and the mismatch in measurement resolution of various covariates, including NDVI and PM2.5 .
@@ 8 Discussion
Method Interval estimation procedure
OLS / RSR (Hodges and Reich, 2010) (unadajusted estimator) Bootstrap.
spline: PLM with spatial effect modeled by thin plate regression spline (Wood, 2003b). Implemented by mgcv (Wood, 2003a). Analytic confidence intervals.
gp: PLM with spatial effect modeled by two-dimensional nearest-neighbors Gaussian process (NNGP) . Implemented by BRISC package (Saha and Datta, 2020). BRISC bootstrap (Saha and Datta, 2018).
gSEM: geo-additive structural equation model (Thaden and Kneib, 2018). Splines fit by mgcv. Bootstrap.
spatial+: spatial+ (Dupont et al., 2020). Splines fit by mgcv. Bootstrap.
svc_mle: Spatially-varying coefficient model (Dambon et al., 2021a). Implemented by varycoef package (Dambon et al., 2021b). Omitted (since resampled points cannot be accommodated by the GP covariance function).
spline_interaction: Spline with spatial interaction of exposure and space (hence, a spatially-varying coefficient model). Implemented by mgcv. Bootstrap.
gp3: Three-dimensional Gaussian process over location and exposure (flexible model). Implemented by RobustGaSP (Gu et al., 2020). Omitted (since resampled points cannot be accommodated by the GP covariance function).
spline3: Three-dimensional spline over location and exposure (flexible model). Implemented by mgcv. Omitted due to computational constraints.
rf: Random forest over location and exposure (flexible model) (Breiman, 2001). Implemented by randomForest package (Liaw and Wiener, 2002). Bootstrap.
BART: Bayesian additive regression tree (BART) over location and exposure (flexible model) (Chipman et al., 2010). Implemented by dbarts (Dorie, 2020) package. Bootstrap.
DML_gp3: gp3 model augmented with spline model over space for exposure (flexible model, DML). Omitted (since resampled points cannot be accommodated by the GP covariance function).
DML_spline: spline3 model augmented with spline model over space for exposure (flexible model, DML). Omitted due to computational constraints.
DML_BART: BART model augmented with spline model over space for exposure (flexible model, DML). Bootstrap.
@@
{adjustwidth}
-3cm-3cm n method bias sd mse coverage $1.000\times 10^{3}$ RSR $7.733\times 10^{-2}$ $1.603\times 10^{-2}$ $6.237\times 10^{-3}$ 0% $1.000\times 10^{3}$ spline $-9.958\times 10^{-5}$ $6.297\times 10^{-3}$ $\mathbf{3.958\times 10^{-5}}$ 96% $1.000\times 10^{3}$ gp $7.932\times 10^{-4}$ $6.509\times 10^{-3}$ $4.292\times 10^{-5}$ 95% $1.000\times 10^{3}$ gSEM $1.362\times 10^{-3}$ $7.017\times 10^{-3}$ $5.099\times 10^{-5}$ 99% $1.000\times 10^{3}$ spatial+ $5.348\times 10^{-4}$ $6.477\times 10^{-3}$ $4.216\times 10^{-5}$ 98% $1.000\times 10^{3}$ svc_mle $3.294\times 10^{-4}$ $6.434\times 10^{-3}$ $4.143\times 10^{-5}$ $1.000\times 10^{3}$ spline_interaction $-9.566\times 10^{-5}$ $6.338\times 10^{-3}$ $4.010\times 10^{-5}$ 99% $1.000\times 10^{3}$ gp3 $-5.282\times 10^{-4}$ $6.478\times 10^{-3}$ $4.216\times 10^{-5}$ $1.000\times 10^{3}$ spline3 $\mathbf{1.538\times 10^{-5}}$ $6.539\times 10^{-3}$ $4.268\times 10^{-5}$ $1.000\times 10^{3}$ rf $-1.886\times 10^{-1}$ $1.370\times 10^{-2}$ $3.575\times 10^{-2}$ 0% $1.000\times 10^{3}$ BART $-2.754\times 10^{-3}$ $1.340\times 10^{-2}$ $1.868\times 10^{-4}$ 100% $1.000\times 10^{3}$ DML_gp3 $-2.531\times 10^{-4}$ $6.690\times 10^{-3}$ $4.473\times 10^{-5}$ $1.000\times 10^{3}$ DML_spline $1.424\times 10^{-4}$ $6.749\times 10^{-3}$ $4.548\times 10^{-5}$ $1.000\times 10^{3}$ DML_BART $-1.811\times 10^{-3}$ $1.350\times 10^{-2}$ $1.850\times 10^{-4}$ 100% $1.000\times 10^{4}$ RSR $7.715\times 10^{-2}$ $5.237\times 10^{-3}$ $5.980\times 10^{-3}$ 0% $1.000\times 10^{4}$ spline $-2.203\times 10^{-4}$ $1.914\times 10^{-3}$ $3.696\times 10^{-6}$ 96% $1.000\times 10^{4}$ gp $-1.842\times 10^{-4}$ $1.978\times 10^{-3}$ $3.932\times 10^{-6}$ 96% $1.000\times 10^{4}$ gSEM $\mathbf{3.427\times 10^{-6}}$ $1.934\times 10^{-3}$ $3.724\times 10^{-6}$ 98% $1.000\times 10^{4}$ spatial+ $-1.298\times 10^{-4}$ $1.912\times 10^{-3}$ $\mathbf{3.657\times 10^{-6}}$ 97% $1.000\times 10^{4}$ spline_interaction $-2.171\times 10^{-4}$ $1.911\times 10^{-3}$ $3.684\times 10^{-6}$ 97% $1.000\times 10^{4}$ spline3 $-1.965\times 10^{-4}$ $1.935\times 10^{-3}$ $3.768\times 10^{-6}$ $1.000\times 10^{4}$ rf $-9.426\times 10^{-2}$ $4.410\times 10^{-3}$ $8.904\times 10^{-3}$ 0% $1.000\times 10^{4}$ BART $-8.981\times 10^{-4}$ $2.987\times 10^{-3}$ $9.694\times 10^{-6}$ 98% $1.000\times 10^{4}$ DML_spline $-1.965\times 10^{-4}$ $1.991\times 10^{-3}$ $3.987\times 10^{-6}$ $1.000\times 10^{4}$ DML_BART $-3.668\times 10^{-4}$ $2.992\times 10^{-3}$ $9.053\times 10^{-6}$ 98%
@@
{adjustwidth}
-3cm-3cm n method bias sd mse coverage $1.000\times 10^{3}$ RSR $2.589\times 10^{-1}$ $1.528\times 10^{-2}$ $6.727\times 10^{-2}$ 0% $1.000\times 10^{3}$ spline $6.863\times 10^{-3}$ $6.502\times 10^{-3}$ $8.930\times 10^{-5}$ 81% $1.000\times 10^{3}$ gp $7.727\times 10^{-3}$ $6.770\times 10^{-3}$ $1.054\times 10^{-4}$ 76% $1.000\times 10^{3}$ gSEM $7.305\times 10^{-2}$ $7.584\times 10^{-3}$ $5.394\times 10^{-3}$ 0% $1.000\times 10^{3}$ spatial+ $1.084\times 10^{-2}$ $6.826\times 10^{-3}$ $1.641\times 10^{-4}$ 73% $1.000\times 10^{3}$ svc_mle $7.387\times 10^{-3}$ $6.855\times 10^{-3}$ $1.015\times 10^{-4}$ $1.000\times 10^{3}$ spline_interaction $6.801\times 10^{-3}$ $6.568\times 10^{-3}$ $8.930\times 10^{-5}$ 96% $1.000\times 10^{3}$ gp3 $2.682\times 10^{-3}$ $6.939\times 10^{-3}$ $5.524\times 10^{-5}$ $1.000\times 10^{3}$ spline3 $6.308\times 10^{-2}$ $4.068\times 10^{-2}$ $5.631\times 10^{-3}$ $1.000\times 10^{3}$ rf $-1.458\times 10^{-1}$ $1.595\times 10^{-2}$ $2.152\times 10^{-2}$ 0% $1.000\times 10^{3}$ BART $7.672\times 10^{-3}$ $1.680\times 10^{-2}$ $3.405\times 10^{-4}$ 100% $1.000\times 10^{3}$ DML_gp3 $\mathbf{1.396\times 10^{-3}}$ $7.065\times 10^{-3}$ $\mathbf{5.177\times 10^{-5}}$ $1.000\times 10^{3}$ DML_spline $2.998\times 10^{-2}$ $3.476\times 10^{-2}$ $2.104\times 10^{-3}$ $1.000\times 10^{3}$ DML_BART $3.432\times 10^{-3}$ $1.685\times 10^{-2}$ $2.950\times 10^{-4}$ 100% $1.000\times 10^{4}$ RSR $2.580\times 10^{-1}$ $4.749\times 10^{-3}$ $6.661\times 10^{-2}$ 0% $1.000\times 10^{4}$ spline $1.630\times 10^{-3}$ $1.947\times 10^{-3}$ $6.433\times 10^{-6}$ 86% $1.000\times 10^{4}$ gp $2.159\times 10^{-3}$ $1.990\times 10^{-3}$ $8.605\times 10^{-6}$ 81% $1.000\times 10^{4}$ gSEM $1.564\times 10^{-2}$ $2.089\times 10^{-3}$ $2.488\times 10^{-4}$ 0% $1.000\times 10^{4}$ spatial+ $5.063\times 10^{-4}$ $1.958\times 10^{-3}$ $\mathbf{4.075\times 10^{-6}}$ 96% $1.000\times 10^{4}$ spline_interaction $1.614\times 10^{-3}$ $1.945\times 10^{-3}$ $6.372\times 10^{-6}$ 88% $1.000\times 10^{4}$ spline3 $2.244\times 10^{-2}$ $6.780\times 10^{-3}$ $5.495\times 10^{-4}$ $1.000\times 10^{4}$ rf $-7.677\times 10^{-2}$ $4.497\times 10^{-3}$ $5.914\times 10^{-3}$ 0% $1.000\times 10^{4}$ BART $2.642\times 10^{-3}$ $3.618\times 10^{-3}$ $2.002\times 10^{-5}$ 97% $1.000\times 10^{4}$ DML_spline $6.870\times 10^{-3}$ $7.080\times 10^{-3}$ $9.712\times 10^{-5}$ $1.000\times 10^{4}$ DML_BART $\mathbf{4.626\times 10^{-4}}$ $3.586\times 10^{-3}$ $1.302\times 10^{-5}$ 100%
@@
{adjustwidth}
-3cm-3cm n method bias sd mse coverage $1.000\times 10^{3}$ RSR $8.670\times 10^{-1}$ $1.173\times 10^{-1}$ $7.655\times 10^{-1}$ 0% $1.000\times 10^{3}$ spline $1.070\times 10^{0}$ $7.224\times 10^{-2}$ $1.150\times 10^{0}$ 0% $1.000\times 10^{3}$ gp $1.063\times 10^{0}$ $7.418\times 10^{-2}$ $1.135\times 10^{0}$ 0% $1.000\times 10^{3}$ gSEM $1.123\times 10^{0}$ $7.587\times 10^{-2}$ $1.268\times 10^{0}$ 0% $1.000\times 10^{3}$ spatial+ $1.090\times 10^{0}$ $7.215\times 10^{-2}$ $1.193\times 10^{0}$ 0% $1.000\times 10^{3}$ svc_mle $2.019\times 10^{-2}$ $3.384\times 10^{-2}$ $1.551\times 10^{-3}$ $1.000\times 10^{3}$ spline_interaction $1.390\times 10^{-2}$ $3.400\times 10^{-2}$ $1.347\times 10^{-3}$ 94% $1.000\times 10^{3}$ gp3 $3.863\times 10^{-3}$ $3.541\times 10^{-2}$ $1.266\times 10^{-3}$ $1.000\times 10^{3}$ spline3 $1.545\times 10^{-1}$ $1.074\times 10^{-1}$ $3.539\times 10^{-2}$ $1.000\times 10^{3}$ rf $-7.837\times 10^{-2}$ $4.106\times 10^{-2}$ $7.824\times 10^{-3}$ 67% $1.000\times 10^{3}$ BART $1.024\times 10^{-1}$ $7.197\times 10^{-2}$ $1.566\times 10^{-2}$ 90% $1.000\times 10^{3}$ DML_gp3 $\mathbf{3.363\times 10^{-3}}$ $3.541\times 10^{-2}$ $\mathbf{1.262\times 10^{-3}}$ $1.000\times 10^{3}$ DML_spline $1.439\times 10^{-1}$ $1.073\times 10^{-1}$ $3.220\times 10^{-2}$ $1.000\times 10^{3}$ DML_BART $9.417\times 10^{-2}$ $7.194\times 10^{-2}$ $1.403\times 10^{-2}$ 92% $1.000\times 10^{4}$ RSR $8.687\times 10^{-1}$ $4.060\times 10^{-2}$ $7.562\times 10^{-1}$ 0% $1.000\times 10^{4}$ spline $1.088\times 10^{0}$ $2.393\times 10^{-2}$ $1.183\times 10^{0}$ 0% $1.000\times 10^{4}$ gp $1.085\times 10^{0}$ $2.416\times 10^{-2}$ $1.178\times 10^{0}$ 0% $1.000\times 10^{4}$ gSEM $1.094\times 10^{0}$ $2.423\times 10^{-2}$ $1.197\times 10^{0}$ 0% $1.000\times 10^{4}$ spatial+ $1.090\times 10^{0}$ $2.389\times 10^{-2}$ $1.189\times 10^{0}$ 0% $1.000\times 10^{4}$ spline_interaction $\mathbf{2.216\times 10^{-3}}$ $1.080\times 10^{-2}$ $\mathbf{1.210\times 10^{-4}}$ 94% $1.000\times 10^{4}$ spline3 $8.317\times 10^{-2}$ $2.091\times 10^{-2}$ $7.353\times 10^{-3}$ $1.000\times 10^{4}$ rf $-4.038\times 10^{-2}$ $1.200\times 10^{-2}$ $1.774\times 10^{-3}$ 16% $1.000\times 10^{4}$ BART $3.066\times 10^{-2}$ $1.514\times 10^{-2}$ $1.168\times 10^{-3}$ 67% $1.000\times 10^{4}$ DML_spline $6.955\times 10^{-2}$ $2.175\times 10^{-2}$ $5.308\times 10^{-3}$ $1.000\times 10^{4}$ DML_BART $2.832\times 10^{-2}$ $1.504\times 10^{-2}$ $1.027\times 10^{-3}$ 74%